# Prym loci of branched double coverings and generalized Andreotti-Mayer loci

### v1

@@ 2. $(2)$
is surjective. This means that $\pi_{1}(\tilde{C}^{(n-2i)},D)$ acts transitively on the fiber $g_{i}^{-1}(D)$, hence $Y_{i}$ is irreducible. Since $0<n-2i\leq 2g+n-1=\tilde{g}$,
$U_{i}=\{D\in\tilde{C}^{(n-2i)}\mid h^{0}(\mathcal{O}_{\tilde{C}}(D))=1,\,\text{$\phi^{*}p\nleq{D}$ for any $p\in{C}$}\}$
is a non-empty open subset of $\tilde{C}^{(n-2i)}$. For $(D_{1},\mathcal{G}_{1}),(D_{2},\mathcal{G}_{2})\in g_{i}^{-1}(U_{i})$, we have
@@
$h^{0}(\mathcal{G}_{1}\otimes\mathcal{G}_{2}^{\vee})=h^{0}(\mathcal{O}_{\tilde{C}}(D_{1})\otimes\phi^{*}(\mathcal{G}_{1}\otimes\mathcal{G}_{2}^{\vee}))=h^{0}(\mathcal{O}_{\tilde{C}}(D_{2}))=1,$
@@
By Proposition 2.3 and Lemma 2.7, we have the following:
@@
Let $(P,[\mathcal{L}])$ be the Prym variety of $[\tilde{C}\overset{\phi}{\rightarrow}{C}]\in\mathcal{R}_{g,2n}$ for $n\geq 1$ and $g\geq 1$.
@@ 1. $(1)$
      $\dim{\operatorname{Bs}{|\mathcal{L}|}}=0$ and #$\#{\operatorname{Bs}{|\mathcal{L}|}}=4^{g}$ for $n=2$,
@@ 3. Quadric hypersurfaces containing projective curves
$\begin{array}[]{ccc}C&\overset{\Phi_{|(\Omega_{C}^{1}\otimes\eta)^{\otimes{i}}|}}{\longrightarrow}&\mathbb{P}(H^{0}((\Omega_{C}^{1}\otimes\eta)^{\otimes{i}}))\\ \Phi\downarrow&\circlearrowleft&\downarrow\rho_{i}^{\vee}\\ \mathbb{P}(H^{0}(\Omega_{C}^{1}\otimes\eta))=\mathbb{P}^{d-1}&\underset{|\mathcal{O}_{\mathbb{P}^{d-1}}(i)|}{\longrightarrow}&\mathbb{P}(\operatorname{Sym}^{i}{H^{0}(\Omega_{C}^{1}\otimes\eta)})=\mathbb{P}(H^{0}(\mathcal{O}_{\mathbb{P}^{d-1}}(i))),\\ \end{array}$
@@ 5. $(5)$
Then $U$ is a non-empty open subset of $C^{g+1}$. We fix $(p_{1},\dots,p_{g+1})\in{U}$, and set $\mathcal{G}_{0}=\mathcal{O}_{C}(p_{1}+\dots+p_{g+1})$ and $\mathcal{G}_{ij}=\mathcal{H}\otimes\mathcal{O}_{C}(p_{1}+\dots+p_{g+1}-p_{i}-p_{j})$ for $1\leq{i}<j\leq{g+1}$. We show that the defining quadratics of $Q_{0,\mathcal{G}}$ and $Q_{0,\mathcal{G}_{ij}}$ form a basis of $\operatorname{Ker}{(\rho_{2})}$. For $E_{0}\neq{E_{\infty}}\in|\mathcal{H}|$, there is a rational function $f$ on $C$ such that $\operatorname{Div}{(f)}=E_{0}-E_{\infty}$. Let $\alpha_{ij}\in{H^{0}(\mathcal{G}_{ij})}$ be a section defining the divisor $E_{\infty}+p_{1}+\dots+p_{g+1}-p_{i}-p_{j}\in|\mathcal{G}_{ij}|$. Then $\alpha^{\prime}_{ij}=f\alpha_{ij}\in{H^{0}(\mathcal{G}_{ij})}$ defines the divisor $E_{0}+p_{1}+\dots+p_{g+1}-p_{i}-p_{j}\in|\mathcal{G}_{ij}|$, and by (3), $\alpha_{ij},\alpha^{\prime}_{ij}$ form a basis of $H^{0}(\mathcal{G}_{ij})$. We set
@@
for $1\leq{k}\leq{g+1}$. Since $\dim{V}=g+1$, by (1) and (5), we have $\bigcap_{k=1}^{g+1}V_{k}=\{0\}$, and $\dim{\bigcap_{k\neq{i}}V_{k}}=1$ for $1\leq{i}\leq{g+1}$. Thus, there is a unique divisor $D_{i}\in|\Omega_{C}^{1}\otimes\eta\otimes\mathcal{H}^{\vee}\otimes\mathcal{O}_{C}(-p_{1}-\dots-p_{g+1}+p_{i})|$. By (1) and (5), we have
@@
Let $D\in|\mathcal{G}_{0}|$ satisfy $\operatorname{supp}{(E_{\infty})}\cap\operatorname{supp}{(D)}=\emptyset$. Then, by (5), we have $V_{D}\cap{V}=\{0\}$ and $\ell_{V}\cap\mathbb{P}(H^{0}(\Omega_{C}^{1}\otimes\eta)/V_{D})=\emptyset$. Since the pencil $|\mathcal{G}_{0}|$ has no base points by (2) and (4), there exist at most two $D\in|\mathcal{G}_{0}|$ such that $\operatorname{supp}{(E_{\infty})}\cap\operatorname{supp}{(D)}\neq\emptyset$. When $p\in\operatorname{supp}{(E_{\infty})}\cap\operatorname{supp}{(D)}$ and $E_{\infty}=p+p^{\prime}$, by (4) we have $p^{\prime}\nleq{D}$. If $V_{D}\subset{V}$, then
@@
and $h^{0}(\Omega_{C}^{1}\otimes\eta\otimes\mathcal{O}_{C}(-D-p-p^{\prime}))\geq 1$. But this contradicts (5), hence $V\nsubseteq{V_{D}}$ and $\ell_{V}\cap\mathbb{P}(H^{0}(\Omega_{C}^{1}\otimes\eta)/V_{D})=\{\Phi(p)\}$. This implies that $\ell_{V}\nsubseteq{Q_{0,\mathcal{G}_{0}}}$. ∎
@@
We prove it by induction on $n$. When $n=4$, the result holds by Lemma 3.2. We assume $n\geq 5$. Let $p_{1},p_{2},p_{3}\in{C\setminus\Sigma}$ be three distinct points. Since the multiplication maps
@@
$=$ $\dim{\operatorname{Ker}{(\rho_{2,p_{1}})}}+\dim{\operatorname{Ker}{(\rho_{2,p_{2}})}}-\dim{\operatorname{Ker}{(\rho_{2,p_{1}+p_{2}})}}=\dim{\operatorname{Ker}{(\rho_{2})}}-1,$
where $\iota_{i}$ denotes the injective homomorphism
$\iota_{i}:\,\operatorname{Sym}^{2}{H^{0}(\Omega_{C}^{1}\otimes\eta\otimes\mathcal{O}_{C}(-p_{i}))}\hookrightarrow\operatorname{Sym}^{2}{H^{0}(\Omega_{C}^{1}\otimes\eta)}.$
@@
$\pi_{i}:\,\mathbb{P}(H^{0}(\Omega_{C}^{1}\otimes\eta))\setminus\{\Phi(p_{i})\}\longrightarrow\mathbb{P}(H^{0}(\Omega_{C}^{1}\otimes\eta\otimes\mathcal{O}_{C}(-p_{i}))).$
@@
Let $(P,[\mathcal{L}])$ be the Prym variety of $[\tilde{C}\overset{\phi}{\rightarrow}{C}]\in\mathcal{R}_{g,2n}$. If $n\geq 4$ and $C$ is an elliptic curve or a hyperelliptic curve, then $\operatorname{Ker}{(\rho_{2})}$ is generated by the defining quadratics of the tangent cones $\mathcal{T}_{\Theta,x}$ for $x\in{S(P,\mathcal{L})}$ and $\Theta\in|\mathcal{L}|$ satisfying $\operatorname{mult}_{x}{\Theta}=2$.
@@
Since $S(P,\mathcal{L})\overset{\simeq}{\rightarrow}S_{2}(P_{\phi},\mathcal{L}_{\phi})$, by Proposition 3.1, it is enough to show that for $E\in{(C\setminus\operatorname{Branch}{(\phi)})^{(n-4)}}$ and $\mathcal{G}\in{K_{2}(E,\eta_{\phi})}$, the quadric $\mathcal{V}_{E,\mathcal{G}}\subset{H^{0}(\Omega_{C}^{1}\otimes\eta_{\phi})^{\vee}}$ is the tangent cone $\mathcal{T}_{W_{\xi},\mathcal{F}}$ for some $\mathcal{F}\in{S_{2}(P_{\phi},\mathcal{L}_{\phi})}$ and $\xi\in\operatorname{Pic}^{0}{(C)}$ satisfying $\operatorname{mult}_{\mathcal{F}}{W_{\xi}}=2$. For $E\in{(C\setminus\operatorname{Branch}{(\phi)})^{(n-4)}}$, there exist $D\in\tilde{C}^{(n-4)}$ such that $\phi_{*}D=E$ and $\phi^{*}p\nleq{D}$ for any $p\in{C}$. Let $\delta\in\operatorname{Pic}^{g+1}{(C)}$ be an invertible sheaf such that $\delta^{\otimes 2}=\Omega_{C}^{1}\otimes\eta_{\phi}\otimes\mathcal{O}_{C}(-E)$. For $\mathcal{G}\in{K_{2}(E,\eta_{\phi})}$, we set $\mathcal{F}=\mathcal{O}_{\tilde{C}}(D)\otimes\phi^{*}\delta$ and $\xi=\mathcal{G}\otimes\delta^{\vee}\in\operatorname{Pic}^{0}{(C)}$. Then, by Lemma 2.5, we have $\operatorname{mult}_{\mathcal{F}}{W_{\xi}}=2$, and $\mathcal{V}_{E,\mathcal{G}}$ is the tangent cone of $W_{\xi}$ at $\mathcal{F}$. ∎
@@ 4. Andreotti-Mayer loci
be the Siegel upper half-space of degree $d=g+n-1$. For the type of the polarization $\Delta=(\underbrace{1,\dots,1}_{n-1},\underbrace{2,\dots,2}_{g})$, we denote the diagonal matrix $\operatorname{diag}{\Delta}\in\operatorname{Mat}_{d}{(\mathbb{Z})}$ by the same notation $\Delta$. We define the action of $\mathbb{Z}^{d}\times\mathbb{Z}^{d}$ on $\mathbb{C}^{d}\times\mathfrak{H}_{d}$ by
$(\mathbb{Z}^{d}\times\mathbb{Z}^{d})\times(\mathbb{C}^{d}\times\mathfrak{H}_{d})\longrightarrow\mathbb{C}^{d}\times\mathfrak{H}_{d};\,(({\boldsymbol{m}}^{\prime},{\boldsymbol{m}}^{\prime\prime}),({\boldsymbol{z}},\tau))\longmapsto({\boldsymbol{z}}+\tau{\boldsymbol{m}}^{\prime\prime}+\Delta{\boldsymbol{m}}^{\prime},\tau),$
@@
$\mathbb{C}^{d}\times\mathbb{C}^{d}\longrightarrow\mathbb{C};\,({\boldsymbol{z}},{\boldsymbol{w}})\longmapsto\operatorname{\mathstrut{}^{\mathit{t}}\!}{{\boldsymbol{z}}}\,(\operatorname{Im}{(\tau)})^{-1}\,\overline{\boldsymbol{w}}.$
@@
Then, by [2, Thenrem 3.2.7. and Remark 8.5.3.],
$\theta_{\boldsymbol{i}}({\boldsymbol{z}},\tau)=\begin{bmatrix}0&\cdots&0&i_{1}&\cdots&i_{g}\\ 0&\cdots&0&0&\cdots&0\end{bmatrix}({\boldsymbol{z}},\tau),\quad\Bigl{(}{\boldsymbol{i}}=\begin{pmatrix}i_{1}\\ \vdots\\ i_{g}\end{pmatrix}\in{I=\Bigl{\{}0,\frac{1}{2}\Bigr{\}}^{g}}\Bigr{)}$
form a basis of $H^{0}(A_{\tau},\mathcal{L}_{\tau})$. We consider a complex analytic space
[…]
Let $\upsilon=u|_{\mathcal{S}}:\mathcal{S}\rightarrow\mathfrak{H}_{d}$ be the restriction of $u:\mathcal{U}_{d}^{\Delta}\rightarrow\mathfrak{H}_{d}$. Then the fiber of $\upsilon$ at $\tau\in\mathfrak{H}_{d}$ is
@@
If $({\boldsymbol{a}},\tau)\in\mathcal{S}$ and ${\boldsymbol{i}}\in{I}$, then $({\boldsymbol{a}}+\tau\begin{pmatrix}{\boldsymbol{0}}\\ {\boldsymbol{i}}\end{pmatrix},\tau)\in\mathcal{S}$.
@@
For ${\boldsymbol{i}},{\boldsymbol{i}}^{\prime}\in{I}$, there exists ${\boldsymbol{i}}^{\prime\prime}\in{I}$ such that ${\boldsymbol{i}}+{\boldsymbol{i}}^{\prime}-{\boldsymbol{i}}^{\prime\prime}\in\mathbb{Z}^{g}$. We set ${\boldsymbol{c}}=\begin{pmatrix}c_{1}\\ \vdots\\ c_{d}\end{pmatrix}=\begin{pmatrix}{\boldsymbol{0}}\\ {\boldsymbol{i}}\end{pmatrix}$. Then we have
$\theta_{{\boldsymbol{i}}^{\prime}}({\boldsymbol{z}}+\tau{\boldsymbol{c}},\tau)=\exp{\Bigl{(}\pi\sqrt{-1}(-\operatorname{\mathstrut{}^{\mathit{t}}\!}{\boldsymbol{c}}\tau{\boldsymbol{c}}-2\operatorname{\mathstrut{}^{\mathit{t}}\!}{\boldsymbol{c}}{\boldsymbol{z}})\Bigr{)}}\theta_{{\boldsymbol{i}}^{\prime\prime}}({\boldsymbol{z}},\tau),$
@@
$\frac{\partial{\theta_{{\boldsymbol{i}}^{\prime}}}}{\partial{z_{j}}}({\boldsymbol{z}}+\tau{\boldsymbol{c}},\tau)\\ =\exp{\Bigl{(}\pi\sqrt{-1}(-\operatorname{\mathstrut{}^{\mathit{t}}\!}{\boldsymbol{c}}\tau{\boldsymbol{c}}-2\operatorname{\mathstrut{}^{\mathit{t}}\!}{\boldsymbol{c}}{\boldsymbol{z}})\Bigr{)}}\bigl{(}-2\pi\sqrt{-1}c_{j}\theta_{{\boldsymbol{i}}^{\prime\prime}}({\boldsymbol{z}},\tau)+\frac{\partial{\theta_{{\boldsymbol{i}}^{\prime\prime}}}}{\partial{z_{j}}}({\boldsymbol{z}},\tau)\bigr{)}$
for $1\leq{j}\leq{d}$. Since $\theta_{{\boldsymbol{i}}^{\prime\prime}}({\boldsymbol{a}},\tau)=0$ and $\frac{\partial{\theta_{{\boldsymbol{i}}^{\prime\prime}}}}{\partial{z_{j}}}({\boldsymbol{a}},\tau)=0$, we have $\theta_{{\boldsymbol{i}}^{\prime}}({\boldsymbol{a}}+\tau{\boldsymbol{c}},\tau)=0$ and $\frac{\partial{\theta_{{\boldsymbol{i}}^{\prime}}}}{\partial{z_{j}}}({\boldsymbol{a}}+\tau{\boldsymbol{c}},\tau)=0$. ∎

### v2

@@ 2. $(2)$
is surjective. This means that $\pi_{1}(\tilde{C}^{(n-2i)},D)$ acts transitively on the fiber $g_{i}^{-1}(D)$, and it follows that $Y_{i}$ is connected. Since $Y_{i}$ is also smooth, it is irreducible. Let $U_{i}$ be the subset of $\tilde{C}^{(n-2i)}$ defined by
$U_{i}=\{D\in\tilde{C}^{(n-2i)}\mid h^{0}(\tilde{C},\mathcal{O}_{\tilde{C}}(D))=1,\,\text{$\phi^{*}p\nleq{D}$ for any $p\in{C}$}\}.$
Since $0<n-2i\leq 2g+n-1=\tilde{g}$, it is a non-empty open subset of $\tilde{C}^{(n-2i)}$. For $(D_{1},\mathcal{G}_{1}),(D_{2},\mathcal{G}_{2})\in g_{i}^{-1}(U_{i})$, we have
@@
$h^{0}(\mathcal{G}_{1}\otimes\mathcal{G}_{2}^{\vee})=h^{0}(\tilde{C},\mathcal{O}_{\tilde{C}}(D_{1})\otimes\phi^{*}(\mathcal{G}_{1}\otimes\mathcal{G}_{2}^{\vee}))=h^{0}(\tilde{C},\mathcal{O}_{\tilde{C}}(D_{2}))=1,$
@@
By Proposition 2.3 and Proposition 2.7, we have the following:
@@
Let $(P,[\mathcal{L}])$ be the Prym variety of $\tilde{C}\overset{\phi}{\rightarrow}{C}]\in\mathcal{R}_{g,2n}$ for $n\geq 1$ and $g\geq 1$.
@@ 1. $(1)$
      $\dim{\operatorname{Bs}{|\mathcal{L}|}}=0$ and ${\operatorname{Bs}{|\mathcal{L}|}}=4^{g}$ for $n=2$,
@@ 3. Quadric hypersurfaces containing projective curves
[…]
@@ 5. $(5)$
For $\mathcal{G}\in\operatorname{Pic}^{g+1}{(C)}$, the condition $\mathcal{G}\in{K(0,\eta)}$ is equivalent to
$h^{0}(\Omega_{C}^{1}\otimes\mathcal{G}^{\vee})=h^{0}(\mathcal{G}\otimes\eta^{\vee})=0.$
Since
$\deg{(\Omega_{C}^{1}\otimes\mathcal{G}^{\vee})}=\deg{(\mathcal{G}\otimes\eta^{\vee})}=g-3<g,$
$K_{2}(0,\eta)=\operatorname{Pic}^{g+1}{(C)}$ for $g\leq 2$, and $K_{2}(0,\eta)$ is a non-empty open subset of $\operatorname{Pic}^{g+1}{(C)}$ with $\dim{(\operatorname{Pic}^{g+1}{(C)}\setminus{K_{2}(0,\eta)})}=g-3$ for $g\geq 3$. Hence the set of $(p_{1},\dots,p_{g+1})\in{C^{g+1}}$ satisfying (2) is a non-empty open subset of $C^{g+1}$. Since the image of the map
$C^{g+1}\longrightarrow\operatorname{Pic}^{g+1}{(C)};\,(p_{1},\dots,p_{g+1})\longmapsto\mathcal{H}\otimes\mathcal{O}_{C}(p_{1}+\dots+p_{g+1}-p_{i}-p_{j})$
has dimension g1g-1 for $1\leq{i}<j\leq{g+1}$, the set of $(p_{1},\dots,p_{g+1})\in{C^{g+1}}$ satisfying (3) is a non-empty open subset of $C^{g+1}$. Since (1), (4) and (5) are also open conditions, $U$ is a non-empty open subset of $C^{g+1}$. We fix $(p_{1},\dots,p_{g+1})\in{U}$, and set $\mathcal{G}_{0}=\mathcal{O}_{C}(p_{1}+\dots+p_{g+1})$ and $\mathcal{G}_{ij}=\mathcal{H}\otimes\mathcal{O}_{C}(p_{1}+\dots+p_{g+1}-p_{i}-p_{j})$ for $1\leq{i}<j\leq{g+1}$. We show that the defining quadratics of $Q_{0,\mathcal{G}}$ and $Q_{0,\mathcal{G}_{ij}}$ form a basis of $\operatorname{Ker}{(\rho_{2})}$. For $E_{0}\neq{E_{\infty}}\in|\mathcal{H}|$, there is a rational function $f$ on $C$ such that $\operatorname{Div}{(f)}=E_{0}-E_{\infty}$. Let $\alpha_{ij}\in{H^{0}(\mathcal{G}_{ij})}$ be a section defining the divisor $E_{\infty}+p_{1}+\dots+p_{g+1}-p_{i}-p_{j}\in|\mathcal{G}_{ij}|$. Then $\alpha^{\prime}_{ij}=f\alpha_{ij}\in{H^{0}(\mathcal{G}_{ij})}$ defines the divisor $E_{0}+p_{1}+\dots+p_{g+1}-p_{i}-p_{j}\in|\mathcal{G}_{ij}|$, and by (3), $\alpha_{ij},\alpha^{\prime}_{ij}$ form a basis of $H^{0}(\mathcal{G}_{ij})$. We set
@@
for $1\leq{k}\leq{g+1}$. Since $\dim{V}=g+1$, by (1) and (5), we have $\bigcap_{k=1}^{g+1}V_{k}=\{0\}$, and $\dim{\bigl(\bigcap_{k\neq{i}}V_{k}\bigr)}=1$ for $1\leq{i}\leq{g+1}$. Thus, there is a unique divisor $D_{i}\in|\Omega_{C}^{1}\otimes\eta\otimes\mathcal{H}^{\vee}\otimes\mathcal{O}_{C}(-p_{1}-\dots-p_{g+1}+p_{i})|$. By (1) and (5), we have
@@
If $D\in|\mathcal{G}_{0}|$ satisfies $\operatorname{supp}{(E_{\infty})}\cap\operatorname{supp}{(D)}=\emptyset$, then, by (5), we have $V_{D}\cap{V}=\{0\}$ and $\ell_{V}\cap\mathbb{P}(H^{0}(\Omega_{C}^{1}\otimes\eta)/V_{D})=\emptyset$. By (2) and (4), since the pencil $|\mathcal{G}_{0}|$ has no base points, there exist at most two $D\in|\mathcal{G}_{0}|$ such that $\operatorname{supp}{(E_{\infty})}\cap\operatorname{supp}{(D)}\neq\emptyset$. When $p\in\operatorname{supp}{(E_{\infty})}\cap\operatorname{supp}{(D)}$ and $E_{\infty}=p+p^{\prime}$, by (4), we have $p^{\prime}\nleq{D}$. If $V_{D}\subset{V}$, then
@@
and $h^{0}(\Omega_{C}^{1}\otimes\eta\otimes\mathcal{O}_{C}(-D-p-p^{\prime}))\geq 1$. But this contradicts (5), hence $V_{D}\nsubseteq{V}$ and $\ell_{V}\cap\mathbb{P}(H^{0}(\Omega_{C}^{1}\otimes\eta)/V_{D})=\{\Phi(p)\}$. This implies that $\ell_{V}\nsubseteq{Q_{0,\mathcal{G}_{0}}}$. ∎
@@
We prove it by induction on $n$. When $n=4$, the result holds by Lemma 3.2. We assume $n\geq 5$. Let $p_{1},p_{2},p_{3}\in{C\setminus\Sigma}$ be three distinct points. We denote by
$\iota_{i}:\,\operatorname{Sym}^{2}{H^{0}(\Omega_{C}^{1}\otimes\eta\otimes\mathcal{O}_{C}(-p_{i}))}\hookrightarrow\operatorname{Sym}^{2}{H^{0}(\Omega_{C}^{1}\otimes\eta)}$
the natural injective homomorphism. Since the multiplication maps
@@
$=$ $\dim{\operatorname{Ker}{(\rho_{2,p_{1}})}}+\dim{\operatorname{Ker}{(\rho_{2,p_{2}})}}-\dim{\operatorname{Ker}{(\rho_{2,p_{1}+p_{2}})}}=\dim{\operatorname{Ker}{(\rho_{2})}}-1.$
@@
$\pi_{i}:\,\mathbb{P}(H^{0}(\Omega_{C}^{1}\otimes\eta))\setminus\{\Phi(p_{i})\}\longrightarrow\mathbb{P}(H^{0}(\Omega_{C}^{1}\otimes\eta\otimes\mathcal{O}_{C}(-p_{i})))$
@@
Let $(P,[\mathcal{L}])$ be the Prym variety of $\tilde{C}\overset{\phi}{\rightarrow}{C}]\in\mathcal{R}_{g,2n}$. If $n\geq 4$ and $C$ is an elliptic curve or a hyperelliptic curve, then $\operatorname{Ker}{(\rho_{2})}$ is generated by the defining quadratics of the tangent cones $\mathcal{T}_{\Theta,x}$ for $x\in{S(P,\mathcal{L})}$ and $\Theta\in|\mathcal{L}|$ satisfying $\operatorname{mult}_{x}{\Theta}=2$.
@@
Since $S(P,\mathcal{L})\overset{\simeq}{\rightarrow}S_{2}(P_{\phi},\mathcal{L}_{\phi})$, by Proposition 3.1, it is enough to show that for $E\in{(C\setminus\operatorname{Branch}{(\phi)})^{(n-4)}}$ and $\mathcal{G}\in{K_{2}(E,\eta_{\phi})}$, the quadric $\mathcal{V}_{E,\mathcal{G}}\subset{H^{0}(\Omega_{C}^{1}\otimes\eta_{\phi})^{\vee}}$ is the tangent cone $\mathcal{T}_{W_{\xi},\mathcal{F}}$ for some $\mathcal{F}\in{S_{2}(P_{\phi},\mathcal{L}_{\phi})}$ and $\xi\in\operatorname{Pic}^{0}{(C)}$ satisfying $\operatorname{mult}_{\mathcal{F}}{W_{\xi}}=2$. For $E\in{(C\setminus\operatorname{Branch}{(\phi)})^{(n-4)}}$, there exist $D\in\tilde{C}^{(n-4)}$ such that $\phi(D)=E$ and $\phi^{*}p\nleq{D}$ for any $p\in{C}$. Let $\delta\in\operatorname{Pic}^{g+1}{(C)}$ be an invertible sheaf such that $\delta^{\otimes 2}=\Omega_{C}^{1}\otimes\eta_{\phi}\otimes\mathcal{O}_{C}(-E)$. For $\mathcal{G}\in{K_{2}(E,\eta_{\phi})}$, we set $\mathcal{F}=\mathcal{O}_{\tilde{C}}(D)\otimes\phi^{*}\delta$ and $\xi=\mathcal{G}\otimes\delta^{\vee}\in\operatorname{Pic}^{0}{(C)}$. Then, by Lemma 2.6, we have $\operatorname{mult}_{\mathcal{F}}{W_{\xi}}=2$, and $\mathcal{V}_{E,\mathcal{G}}$ is the tangent cone of $W_{\xi}$ at $\mathcal{F}$. ∎
@@ 4. Andreotti-Mayer loci
be the Siegel upper half-space of degree d=g+n1d=g+n-1. For the type of the polarization $\Delta=(\underbrace{1,\dots,1}_{n-1},\underbrace{2,\dots,2}_{g})$, we denote the diagonal matrix $\operatorname{diag}{\Delta}\in\operatorname{Mat}_{d}{(\mathbb{Z})}$ by the same notation $\Delta$. We define the action of $\mathbb{Z}^{d}\times\mathbb{Z}^{d}$ on $\mathbb{C}^{d}\times\mathfrak{H}_{d}$ by
$(\mathbb{Z}^{d}\times\mathbb{Z}^{d})\times(\mathbb{C}^{d}\times\mathfrak{H}_{d})\longrightarrow\mathbb{C}^{d}\times\mathfrak{H}_{d};\,(({\boldsymbol{m}}^{\prime},{\boldsymbol{m}}^{\prime\prime}),({\boldsymbol{z}},\tau))\longmapsto({\boldsymbol{z}}+\tau{\boldsymbol{m}}^{\prime}+\Delta{\boldsymbol{m}}^{\prime\prime},\tau),$
@@
$H:\,\mathbb{C}^{d}\times\mathbb{C}^{d}\longrightarrow\mathbb{C};\,({\boldsymbol{z}},{\boldsymbol{w}})\longmapsto\operatorname{\mathstrut{}^{\mathit{t}}\!}{{\boldsymbol{z}}}\,(\operatorname{Im}{(\tau)})^{-1}\,\overline{\boldsymbol{w}}$
and the semi-character
$\chi_{0}:\,\tau\mathbb{Z}^{d}+\Delta\mathbb{Z}^{d}\longrightarrow\mathbb{C}^{*};\,\tau{\boldsymbol{m}}^{\prime}+\Delta{\boldsymbol{m}}^{\prime\prime}\longmapsto\exp{\bigl(\pi\sqrt{-1}\,\operatorname{\mathstrut{}^{\mathit{t}}\!}{\boldsymbol{m}}^{\prime}\Delta{\boldsymbol{m}}^{\prime\prime}\bigr)}.$
@@
Then, by [2, Theorem 3.2.7. and Remark 8.5.3.],
$\theta_{\boldsymbol{i}}({\boldsymbol{z}},\tau)=\theta\begin{bmatrix}0&\cdots&0&i_{1}&\cdots&i_{g}\\ 0&\cdots&0&0&\cdots&0\end{bmatrix}({\boldsymbol{z}},\tau),\quad\Bigl({\boldsymbol{i}}=\begin{pmatrix}i_{1}\\ \vdots\\ i_{g}\end{pmatrix}\in{I=\Bigl\{0,\frac{1}{2}\Bigr\}^{g}}\Bigr)$
form a basis of $H^{0}(A_{\tau},\mathcal{L}_{\tau})$. We consider the complex analytic space
$\mathcal{S}=\bigl\{({\boldsymbol{z}},\tau)\in\mathcal{U}_{d}^{\Delta}\mid\theta_{\boldsymbol{i}}({\boldsymbol{z}},\tau)=0,\,\frac{\partial\theta_{\boldsymbol{i}}}{\partial{z_{j}}}({\boldsymbol{z}},\tau)=0\quad({\boldsymbol{i}}\in{I},\,1\leq{j}\leq{d})\bigr\},$
where $z_{j}$ denotes the coordinate of ${\boldsymbol{z}}=\begin{pmatrix}z_{1}\\ \vdots\\ z_{d}\end{pmatrix}\in\mathbb{C}^{d}$. Let $\upsilon=u|_{\mathcal{S}}:\mathcal{S}\rightarrow\mathfrak{H}_{d}$ be the restriction of $u:\mathcal{U}_{d}^{\Delta}\rightarrow\mathfrak{H}_{d}$. Then the fiber of $\upsilon$ at $\tau\in\mathfrak{H}_{d}$ is
@@
If $({\boldsymbol{a}},\tau)\in\mathcal{S}$, then $\bigl({\boldsymbol{a}}+\tau\begin{pmatrix}{\boldsymbol{0}}\\ {\boldsymbol{i}}\end{pmatrix}+\Delta\begin{pmatrix}{\boldsymbol{0}}\\ {\boldsymbol{j}}\end{pmatrix},\tau\bigr)\in\mathcal{S}$ for ${\boldsymbol{i}},{\boldsymbol{j}}\in{I}$.
@@
For ${\boldsymbol{i}},{\boldsymbol{i}}^{\prime}\in{I}$, there exists ${\boldsymbol{i}}^{\prime\prime}\in{I}$ such that ${\boldsymbol{i}}+{\boldsymbol{i}}^{\prime}-{\boldsymbol{i}}^{\prime\prime}\in\mathbb{Z}^{g}$. We set ${\boldsymbol{c}}=\begin{pmatrix}c_{1}\\ \vdots\\ c_{d}\end{pmatrix}=\begin{pmatrix}{\boldsymbol{0}}\\ {\boldsymbol{i}}\end{pmatrix}\in\frac{1}{2}\mathbb{Z}^{d}$ and ${\boldsymbol{d}}=\Delta\begin{pmatrix}{\boldsymbol{0}}\\ {\boldsymbol{j}}\end{pmatrix}\in\mathbb{Z}^{d}$. Then we have
$\theta_{{\boldsymbol{i}}^{\prime}}({\boldsymbol{z}}+\tau{\boldsymbol{c}}+{\boldsymbol{d}},\tau)=\exp{\Bigl(\pi\sqrt{-1}(-\operatorname{\mathstrut{}^{\mathit{t}}\!}{\boldsymbol{c}}\tau{\boldsymbol{c}}-2\operatorname{\mathstrut{}^{\mathit{t}}\!}{\boldsymbol{c}}{\boldsymbol{z}}+4\operatorname{\mathstrut{}^{\mathit{t}}\!}{\boldsymbol{i}}^{\prime}{\boldsymbol{j}})\Bigr)}\theta_{{\boldsymbol{i}}^{\prime\prime}}({\boldsymbol{z}},\tau),$
@@
$\frac{\partial{\theta_{{\boldsymbol{i}}^{\prime}}}}{\partial{z_{j}}}({\boldsymbol{z}}+\tau{\boldsymbol{c}}+{\boldsymbol{d}},\tau)\\ =\exp{\Bigl(\pi\sqrt{-1}(-\operatorname{\mathstrut{}^{\mathit{t}}\!}{\boldsymbol{c}}\tau{\boldsymbol{c}}-2\operatorname{\mathstrut{}^{\mathit{t}}\!}{\boldsymbol{c}}{\boldsymbol{z}}+4\operatorname{\mathstrut{}^{\mathit{t}}\!}{\boldsymbol{i}}^{\prime}{\boldsymbol{j}})\Bigr)}\bigl(-2\pi\sqrt{-1}\,c_{j}\theta_{{\boldsymbol{i}}^{\prime\prime}}({\boldsymbol{z}},\tau)+\frac{\partial{\theta_{{\boldsymbol{i}}^{\prime\prime}}}}{\partial{z_{j}}}({\boldsymbol{z}},\tau)\bigr)$
for $1\leq{j}\leq{d}$. Since $\theta_{{\boldsymbol{i}}^{\prime\prime}}({\boldsymbol{a}},\tau)=0$ and $\frac{\partial{\theta_{{\boldsymbol{i}}^{\prime\prime}}}}{\partial{z_{j}}}({\boldsymbol{a}},\tau)=0$, we have $\theta_{{\boldsymbol{i}}^{\prime}}({\boldsymbol{a}}+\tau{\boldsymbol{c}}+{\boldsymbol{d}},\tau)=0$ and $\frac{\partial{\theta_{{\boldsymbol{i}}^{\prime}}}}{\partial{z_{j}}}({\boldsymbol{a}}+\tau{\boldsymbol{c}}+{\boldsymbol{d}},\tau)=0$. ∎